OpenStudy (aroub):

Triangle PQR is equilateral with altitude of length 7sqrt3. What is the perimeter of triangle PQR?

OpenStudy (anonymous):

Come on aroub you can do it!

OpenStudy (accessdenied):

i think if you draw an equilateral triangle and place the altitude in, you can kind of see what' going on better...

OpenStudy (cwrw238):

one side = 2*7 = 14 perimeter = 42

OpenStudy (aroub):

What's the perimeter?

OpenStudy (anonymous):

Actually, Altitude of an equilateral triangle is \[ \frac{\sqrt{3}}2 \times side \]

OpenStudy (cwrw238):

altude forms a 60-30-90 triangle so sides are in ratio 2:1: sqrt3

OpenStudy (cwrw238):

in this case side of equilateral triangle = = 7*2 = 14

OpenStudy (aroub):

 I mean the formula!

OpenStudy (aroub):

P= PQ+QR+PR ?

OpenStudy (accessdenied):

yes, that would be perimeter around the triangle

OpenStudy (anonymous):

Yes, also note that in equilateral triangle the sides are equal.